### v1

# Unlearning Isn’t Deletion: Investigating Reversibility of Machine Unlearning in LLMs

Abstract
Unlearning in large language models (LLMs) is intended to remove the influence of specific data, yet current evaluations rely heavily on token-level metrics such as accuracy and perplexity. We show that these metrics can be misleading: models often appear to forget, but their original behavior can be rapidly restored with minimal fine-tuning, revealing that unlearning may obscure information rather than erase it. To diagnose this phenomenon, we introduce a representation-level evaluation framework using PCA-based similarity and shift, centered kernel alignment, and Fisher information. Applying this toolkit across six unlearning methods, three domains (text, code, math), and two open-source LLMs, we uncover a critical distinction between reversible and irreversible forgetting. In reversible cases, models suffer token-level collapse yet retain latent features; in irreversible cases, deeper representational damage occurs. We further provide a theoretical account linking shallow weight perturbations near output layers to misleading unlearning signals, and show that reversibility is modulated by task type and hyperparameters. Our findings reveal a fundamental gap in current evaluation practices and establish a new diagnostic foundation for trustworthy unlearning in LLMs. We provide a unified toolkit for analyzing LLM representation changes under unlearning and relearning: https://github.com/XiaoyuXU1/Representational_Analysis_Tools.git.
1 Introduction
Large language models (LLMs), trained on massive corpora, have achieved remarkable success across diverse tasks, yet their capacity to memorize training snippets poses acute ethical, legal, and security risks. Memorization can unintentionally disclose sensitive, harmful, or copyrighted text [25; 13; 30], conflicting with emerging regulations, such as the EU’s Right to be Forgotten [9].
Machine unlearning seeks to mitigate this threat by making a model act as though specified data were never seen [2]. Numerous methods have been proposed for LLMs [33; 12; 6; 26; 18; 17; 19; 20; 31], their success usually judged by token-level metrics, such as accuracy or perplexity.
However, a pivotal question remains largely unexplored: Does LLM unlearning truly erase information, or merely suppress it, poised to “resurface” at the slightest nudge?
Empirically, many unlearning methods may just appear effective: after unlearning, a model shows near-zero accuracy or high perplexity on the forget set, but a brief fine-tuning step (even on unrelated data) can quickly restore its original behaviour (see Figure 1). This exposes a huge gap between surface-level metrics and the model’s internal state [21; 23], casting doubt on compliance and safety claims. If information is recoverable via simple fine-tuning (or relearning), can we truly claim that it has been “forgotten?” What looks like memory loss may, in fact, be a shallow perturbation.
In this work, we perform the first systematic analysis of reversibility of LLM unlearning, covering both single-shot and continual settings. The continual setting allows multiple unlearning requests over time, which we believe is a more prominent and realistic setting where the deployment of such unlearning strategies are often facing a dynamic environment. We show that standard token-level metrics prove insufficient—they can collapse even when the underlying representations remain intact. To probe deeper, we introduce a diagnostic toolkit for representational analysis, featuring PCA subspace similarity and shift [36], centered kernel alignment (CKA) [15], and Fisher information [3]. Our toolkit uncovers two distinct regimes of unlearning:
1) Reversible (catastrophic) forgetting: performance collapses, but feature subspaces are largely preserved, enabling rapid recovery, and 2) Irreversible (catastrophic) forgetting: collapse coincides with substantial representational drift, making recovery difficult/impossible. Surprisingly, both yield similar results under token-level metrics, underscoring the need for deeper representational analysis.
We further show that the transition between reversible and irreversible forgetting depends not only on the volume of unlearning requests but also on hyperparameters such as learning rate. Modest weight perturbations—especially near the output layer—can lead to token-level distortions without altering feature geometry, making “forgotten” knowledge easily recoverable.
Hence, evaluating the effectiveness of LLM unlearning must go beyond superficial token-level metrics (e.g., forget accuracy declines). In safety- and privacy-critical settings, unlearning should be judged by its ability to achieve genuine erasure rather than simply representational collapse.
Figure 1: (a) Token-level accuracy and CKA subspaces of reversible (top) vs. irreversible (bottom) catastrophic forgetting due to continual unlearning then relearning, (b) Our four diagnostic tools
Contributions. We summarize our main contributions as follows:
  • We present the first systematic study of reversibility in both single and continual LLM unlearning, using a feature-space toolkit, including PCA similarity, PCA shift, CKA, and Fisher information. Our analysis distinguishes reversible from irreversible (catastrophic) forgetting.
  • We conduct extensive experiments with six unlearning methods (GA [33], NPO [35], and RLabel, with their variants) across three datasets (arXiv papers, GitHub code [33], and NuminaMath-1.5 [16]) on Yi-6B [34] and Qwen-2.5-7B [32]. Our results show that standard token-level metrics (e.g., accuracy, perplexity, MIA susceptibility [27]) fail to capture true forgetting behavior.
  • We theoretically analyze weight perturbations to explain how widespread vs. localized parameter changes relate to (ir)reversible forgetting. Small perturbations near the logits can distort token-level metrics despite intact features, hence leading to misleading assessments.
  • Based on our findings and extra preliminary results, we propose several future directions, including using unlearning as a complementary form of data augmentation, and designing more robust unlearning algorithms that achieve genuine forgetting while avoiding representational collapse.
2 Preliminaries
LLM unlearning seeks to enhance privacy, improve safety, and reduce bias [33; 12; 26; 18; 17; 19]. Most work adopts the single-unlearning paradigm: given a training corpus $\mathcal{D}$ and a designated forget set $\mathcal{D}_{f}\subseteq\mathcal{D}$, a model $\mathcal{M}$ is first trained on $\mathcal{D}$ with algorithm $\mathcal{A}$. An unlearning procedure $\mathcal{U}$ then transforms $\mathcal{M}$ into an unlearned model $\mathcal{M}_{f}$ that should behave as if it had never encountered $\mathcal{D}_{f}$.
Ideally, $\mathcal{U}$ should produce a model statistically indistinguishable from one retrained on the retain set $\mathcal{D}_{r}=\mathcal{D}\setminus\mathcal{D}_{f}$: $\mathcal{M}_{f}=\mathcal{U}\!\bigl{(}\mathcal{M},\mathcal{D}_{f}\bigr{)}\approx\mathcal{M}_{r}=\mathcal{A}\!\bigl{(}\mathcal{M},\mathcal{D}_{r}\bigr{)}$. While current methods strike a good forget-utility balance in controlled settings [1; 8] (e.g., a fixed $\mathcal{D}_{f}$ or just a single removal request), they seldom address the practical need for continual unlearning, i.e., data owners can submit removal requests sequentially over time [1]. Let the successive forget sets be $\mathcal{D}_{f}^{(1)},\mathcal{D}_{f}^{(2)},\ldots,\mathcal{D}_{f}^{(t)}$ (whose union is $\mathcal{D}_{f}$); the retain set after $t$ rounds is $\mathcal{D}_{r}^{(t)}$. The model is then updated recursively: $\mathcal{M}_{f}^{(t+1)}=\mathcal{U}(\mathcal{M}_{f}^{(t)},\mathcal{D}_{f}^{(t+1)})$, which should be similar to $\mathcal{M}_{r}=\mathcal{A}\!\bigl{(}\mathcal{M},\mathcal{D}_{r}^{(t+1)}\bigr{)}$ at any time $t$.
Retraining LLMs is prohibitively costly, so most studies rely on empirical proxies rather than formal statistically-indistinguishable guarantees [6; 24; 18; 7]. Evaluations track forget quality on the forget set and utility on the retain set, aiming to preserve both [24]. While single unlearning often yields modest declines, it is fragile: brief fine-tuning—even on benign, unrelated data—swiftly revives the “forgotten” knowledge [1; 23; 21]. The issue worsens under continual unlearning, where each round begins from an already degraded model, ultimately triggering catastrophic forgetting—a wholesale performance collapse [1; 28]. Prior work notes this risk but does not examine its root causes.
We hypothesize that collapse does not imply true erasure; the knowledge may remain latent in the feature space. This insight leads us to distinguish two regimes of (catastrophic) forgetting:
Definition 1 (Reversible (Catastrophic) Forgetting).
Let $\theta_{0}$ denote the initial model parameters, $\mathcal{D}_{f}$ the forget set, and $\mathcal{T}$ an evaluation task with metric $E(\cdot,\mathcal{T})$. Unlearning $\mathcal{D}_{f}$ transforms the model to $\theta_{u}$. If subsequent relearning on $\mathcal{D}_{f}$ (or an equivalent reconstruction set) produces parameters $\theta_{r}$ s.t.
$E(\theta_{u},\mathcal{T})\;\ll\;E(\theta_{r},\mathcal{T})\;\approx\;E(\theta_{0},\mathcal{T}),$
the temporary performance collapse is fully reversible; we call it reversible catastrophic forgetting. When the initial degradation is modest (e.g., single unlearning), we simply call it reversible forgetting.
Definition 2 (Irreversible (Catastrophic) Forgetting).
Using the notations of Definition 1, if
$E(\theta_{u},\mathcal{T})\;\approx\;E(\theta_{r},\mathcal{T})\;\ll\;E(\theta_{0},\mathcal{T}),$
we observe irreversible catastrophic forgetting: the collapse (i.e., weight perturbation) is irreversible. We refine this to irreversible forgetting, which further requires that the irreversible degradation be restricted to the forget set; performance on the retain set and on unrelated data must remain near their original levels. This condition distinguishes targeted erasure from global model failure or collapse.
To distinguish our setting from a full retrain model, where reversibility is trivially achievable, we introduce the following restriction on the relearning phase.
Relearning Restriction. After unlearning, we briefly fine-tune $\theta_{u}$ on a small relearning set—either the cumulative forget set $\mathcal{D}_{f}=\bigcup_{t}\mathcal{D}_{f}^{(t)}$, its a distribution-similar retain set $\mathcal{D}_{r}^{(t)}$, or an unrelated out-of-distribution corpus—to obtain $\theta_{r}$ without ever revisiting the full pre-training data.
3 Token-Level Evaluation
3.1 Experiment setup
Models and Datasets. We conduct experiments on two open-source models, Yi-6B [34] and Qwen-2.5-7B [32]. To ensure the generality of our findings, we use two dataset types: i) simple tasks—arXiv abstracts and GitHub code from [33], and ii) a complex task—NuminaMath-1.5, a recent benchmark for mathematical reasoning [16]. All experiments were run on NVIDIA H100 GPUs.
Unlearning algorithms We compare six canonical methods grouped into three families.
1) Gradient-Ascent (GA) family. The unified goal is $\mathcal{L}\;=\;\mathcal{L}_{\text{forget}}\!\bigl{(}\mathcal{D}_{f}\bigr{)}+\;\lambda\,\mathcal{L}_{\text{retain}}\!\bigl{(}\mathcal{D}_{r}\bigr{)},$ where $\mathcal{L}_{\mathrm{forget}}$ maximizes the loss on the forget set via GA, $\mathcal{L}_{\mathrm{retain}}$ (optional) preserves utility on the retain set, and $\lambda>0$ balances the two. Choices for $\mathcal{L}_{\mathrm{retain}}$ give three variants: i) GA ($\mathcal{L}_{\mathrm{retain}}=0$), ii) GA+GD (standard cross-entropy on $\mathcal{D}{r}$), and iii) GA+KL (KL divergence to the reference model on $\mathcal{D}{r}$[33].
2) Negative Preference Optimization (NPO) family. GA is replaced by an NPO loss that penalizes agreement with the forget set [35]: $\mathcal{L}\;=\;\mathcal{L}_{\text{NPO}}\!\bigl{(}\mathcal{D}_{f}\bigr{)}+\;\lambda\,\mathcal{L}_{\text{retain}}\!\bigl{(}\mathcal{D}_{r}\bigr{)},$ Variants mirror those above: NPO ($\mathcal{L}_{\mathrm{retain}}=0$) and NPO+KL (retain-set KL regularization).
3) Random Label (RLabel). To mimic a model that never saw $\mathcal{D}_{f}$, true labels are replaced with random ones: $\mathcal{L}\;=\;\mathcal{L}_{\text{RLabel}}\!\bigl{(}\mathcal{D}_{f}\bigr{)},$ inducing near-uniform predictions without GA/negative rewards [33].
Unlearning Scenario We consider two standard settings: i) Single unlearning: A trained model $\mathcal{M}$ receives exactly one request to remove $\mathcal{D}_{f}\subset\mathcal{D}$, producing $\mathcal{M}_{f}=\mathcal{U}(\mathcal{M},\mathcal{D}_{f})$, and ii) Continual unlearning: The model processes a stream of requests ${\mathcal{D}_{f}^{(1)},\dots,\mathcal{D}_{f}^{(k)}}$, updated iteratively by $\mathcal{M}^{(i+1)}=\mathcal{U}(\mathcal{M}^{(t)},\mathcal{D}_{f}^{(i+1)})$ with $\mathcal{M}^{(0)}=\mathcal{M}$ and $\bigcup_{i=1}^{n}\mathcal{D}_{f}^{(t)}=\mathcal{D}_{f}$. This mirrors real-world, incremental removal demands while maintaining parity with the single-step budget.
For the simple tasks, we benchmark all six algorithms: GA, GA + GD, GA + KL, RLabel, NPO, and NPO+KL. For the complex one, lacking a clearly defined retain set, we use GA, NPO, and RLabel.
Evaluation Metrics For single unlearning (simple tasks only), we report: forget-set accuracy (F.Acc), retain-set accuracy (R.Acc), and privacy leakage via min-$k\%$-prob MIA AUC [27].
[…]
Relearning setting. To gauge how readily forgotten knowledge re-emerges, each unlearning run is followed by a controlled relearning phase. Single unlearning: We fine-tune on the whole forget set $\mathcal{D}_{f}$ once, producing a single-step relearned model. Continual unlearning: For settings that trigger catastrophic collapse, we fine-tune with three cases: i) the cumulative forget set $\bigcup_{i}\mathcal{D}_{f}^{(t)}$, ii) the corresponding retain set $\mathcal{D}_{r}^{(t)}$, and iii) an unrelated auxiliary corpus. These progressively relax assumptions about access to the forgotten content, revealing the recovery potential in each case.
Hyperparameter Configuration. To comprehensively evaluate the effects of unlearning, we design multiple hyperparameter configurations that vary both the learning rate and the number of unlearning requests. For single unlearning we sweep the learning rate over $LR\in\{3,4,5\}\!\times\!10^{-6}$ while fixing the request count to $N=1$. For continual unlearning we vary both knobs: on the simple task (Yi-6B) we test $LR\in\{3,5\}\!\times\!10^{-6}\cup\{3\times 10^{-5}\}$ with $N\in\{6\to 100\}$; on the complex task (Qwen-2.5-7B) we use $LR\in\{3,5\}\!\times\!10^{-6}$ and $3\times 10^{-5}$ together with $N\in\{6\to 100\}$. All runs adopt the optimizer settings of [1]: AdamW [22] ($\beta_{1}=0.9$, $\beta_{2}=0.95$, $\varepsilon=10^{-8}$), a cosine schedule with 10% warm-up followed by decay to 10% of peak, weight decay 0.1, and gradient clipping at 1.0.
Table 1: Yi-6B: MIA / F.Acc / R.Acc (%) simple task using three LRs under single unlearning
Phase Method LR=3$\times$10-6 LR=4$\times$10-6 LR=5$\times$10-6
MIA F.Acc R.Acc MIA F.Acc R.Acc MIA F.Acc R.Acc
Original 70.9 78.9 65.5 70.9 78.9 65.5 70.9 78.9 65.5
Unlearn GA 45.5 65.4 54.0 43.8 62.4 52.3 41.2 60.3 50.9
GA+GD 65.4 75.1 64.6 58.2 73.8 65.8 55.3 68.5 63.5
GA+KL 48.9 71.0 58.5 47.6 70.6 58.1 44.8 68.4 55.4
NPO 67.2 76.2 64.7 65.2 75.8 62.8 62.2 75.2 62.7
NPO+KL 66.5 76.3 64.8 67.2 76.4 63.2 64.5 75.6 61.2
RLabel 69.6 77.7 64.7 69.2 76.5 64.5 68.7 75.4 63.3
Relearn GA 67.2 76.6 65.2 68.6 77.6 62.8 67.6 76.9 65.5
GA+GD 68.6 77.0 65.3 68.8 76.9 65.3 68.8 77.2 65.3
GA+KL 67.9 77.6 65.3 68.3 75.5 65.2 67.7 77.2 65.2
NPO 68.2 77.1 65.3 68.2 77.2 65.2 68.3 77.0 65.1
NPO+KL 68.9 77.1 65.3 67.9 76.3 63.0 68.6 76.9 65.2
RLabel 68.3 78.8 65.6 68.9 76.4 65.3 68.8 78.9 65.2
3.2 Token-Level Evaluation Results
We report quantitative results on both single unlearning and continual unlearning settings using Yi-6B and Qwen-2.5-7B across multiple configurations (Tables 13). For completeness, detailed experimental results and additional settings are provided in Appendix A.4.
Single Unlearning. For Yi-6B under single unlearning, Table 1 shows that all six unlearning methods reduce MIA and F.Acc, indicating some level of unlearning. The changes on retain set are modest: R.Acc only drops 2–5% for most methods, and MIA decreases by less than 30 points in most cases. Importantly, the relearned models often recover original performance—both GA+KL and RLabel restore R.Acc to near 65.0%, and F.Acc rebounds above 77%. These results suggest that under single unlearning, most methods achieve seemingly successful forgetting at the token level, but as we show in Section 4.2 to Section 4.2, the underlying representation changes are minimal—indicating the phenomenon of reversible forgetting.
Continual Unlearning. By examining post-relearning recoverability in Table 2 and Table 3, we identify two distinct forms of catastrophic forgetting. When the model regains both utility (e.g., F.Acc, R.Acc) and privacy (e.g., MIA AUC) to levels near or exceeding the original after relearning, we classify the behavior as reversible catastrophic forgetting. This suggests that the underlying representational structure remains intact, enabling efficient recovery via lightweight retraining. Such reversibility is consistently observed in methods like NPO and NPO+KL, particularly under low learning rates or small removal batches.
Conversely, when relearning fails to restore utility—reflected in persistently low F.Acc and R.Acc despite partial MIA recovery—we categorize it as irreversible catastrophic forgetting. This scenario frequently arises with methods like GA and RLabel under aggressive hyperparameters (e.g., $\mathrm{LR}=3\times 10^{-5}$), where damage accumulates across layers and results in irreversible representational collapse. Importantly, MIA AUC alone can be misleading. Models may exhibit near-complete privacy recovery while remaining functionally impaired.
Table 2: Yi-6B: MIA / F.Acc / R.Acc (%) for simple task under four unlearning settings
Phase Method LR=3$\times$$10^{-5}$, N=100 LR=5$\times$$10^{-6}$, N=100 LR=3$\times$$10^{-6}$, N=100 LR=3$\times$$10^{-5}$, N=6
MIA F.Acc R.Acc MIA F.Acc R.Acc MIA F.Acc R.Acc MIA F.Acc R.Acc
Original —— 70.8 78.9 65.5 70.8 78.9 65.5 70.8 78.9 65.5 70.8 78.9 65.5
Unlearn GA 26.1 0.0 0.0 23.2 9.1 6.2 25.2 16.8 14.4 29.6 36.3 36.1
GA+GD 16.8 9.7 2.3 28.7 3.6 3.1 69.4 78.8 65.5 66.9 77.0 64.0
GA+KL 17.8 9.0 6.2 27.3 9.1 6.2 18.9 3.8 3.2 29.5 52.9 41.5
NPO 60.1 37.8 37.9 50.6 51.0 52.3 68.4 78.3 64.1 68.7 71.6 59.4
NPO+KL 59.0 64.3 55.9 65.4 77.6 64.3 66.7 78.8 65.5 67.9 67.6 56.1
RLabel 65.1 0.0 0.0 63.6 0.1 0.4 61.4 0.4 0.7 62.7 72.7 61.1
Relearn GA 74.5 2.1 1.8 68.0 80.0 65.0 68.6 80.8 65.2 68.2 70.5 58.7
GA+GD 68.1 2.2 2.6 69.8 81.2 65.1 70.0 81.8 65.5 67.0 61.6 54.4
GA+KL 70.7 1.7 1.6 68.3 81.1 64.8 70.7 81.0 63.2 65.0 66.6 56.2
NPO 70.0 57.0 45.6 68.0 82.7 65.5 69.9 81.2 65.4 68.4 71.2 59.4
NPO+KL 67.7 60.7 54.2 69.5 83.8 65.6 69.9 83.8 65.4 69.0 67.6 56.1
RLabel 69.5 4.3 2.8 70.4 80.8 65.3 70.0 80.5 65.3 65.2 72.7 61.1
Table 3: Qwen-2.5-7B: MIA / MATH / GSM8K Accuracy (%) for complex task under four settings
Phase Method LR=$3\times 10^{-5}$, $N$=6 LR=$3\times 10^{-6}$, $N$=6 LR=$5\times 10^{-6}$, $N$=6 LR=$5\times 10^{-6}$, $N$=100
MIA MATH GSM8K MIA MATH GSM8K MIA MATH GSM8K MIA MATH GSM8K
Original —– 99.3 9.0 80.1 99.3 9.0 80.1 99.3 9.0 80.1 99.3 9.0 80.1
Unlearn GA 5.9 0.0 0.0 0.9 0.0 0.0 3.8 0.0 0.0 5.5 0.0 0.0
NPO 95.9 0.0 0.2 97.4 21.5 74.1 67.4 24.1 71.8 94.7 0.0 0.4
RLabel 35.5 0.0 0.0 69.6 0.0 1.5 11.2 0.0 0.0 2.9 0.0 0.0
Relearn GA 97.6 0.0 1.1 99.3 5.1 83.2 99.4 9.3 77.8 99.2 0.0 0.0
NPO 95.8 0.0 0.0 99.4 4.7 82.6 99.4 16.5 75.7 99.2 0.0 0.0
RLabel 99.5 0.0 0.0 99.3 5.3 83.3 99.3 10.0 77.2 99.6 0.0 0.0
4 A Unified Representational Analysis
4.1 Representational Analysis Tools
We monitor representational drift with four layer-wise diagnostics: PCA Similarity, PCA Shift, CKA, and the diagonal Fisher Information Matrix (FIM), summarized in Figure 1(b); implementation details (including their precise definitions) are deferred to Appendix A.3.
PCA Similarity & Shift. For each layer $i$, we collect activations $\mathbf{H}_{i}^{\text{orig}}$, $\mathbf{H}_{i}^{\text{unl}}$, and $\mathbf{H}_{i}^{\text{rel}}$ on a probe set $\mathcal{X}$. Let $\mathbf{c}_{i,1}^{(*)}$ and $p_{i,1}^{(*)}$ be he first principal direction and its mean projection for state (){$(*)\!\in\!\{\text{orig},\text{unl},\text{rel}\}$. The cosine between $\mathbf{c}_{i,1}^{\text{orig}}$ and $\mathbf{c}_{i,1}^{(*)}$ yields PCA similarity; the signed difference $p_{i,1}^{(*)}-p_{i,1}^{\text{orig}}$ gives PCA shift. Small angles and shifts indicate stable features; otherwise, catastrophic forgetting [36].
Centered Kernel Alignment (CKA). With centered activation matrices $X_{i}^{\text{orig}}$ and $X_{i}^{(*)}$, we compute ${\rm CKA}(X_{i}^{\text{orig}},X_{i}^{(*)})\!\in\![0,1]$; values $\approx 1$ mean nearly identical subspaces, those $\approx 0$ are orthogonal.
Fisher information. We estimate the diagonal empirical FIM by averaging squared gradients over $\mathcal{X}$. Comparing $\mathrm{FIM}^{\text{orig}}$, $\mathrm{FIM}^{\text{unl}}$, and $\mathrm{FIM}^{\text{rel}}$ reveals how unlearning flattens the loss landscape and whether relearning restores parameter importance [14; 11].
We compute all diagnostics not only on the forget set but also on the retain set and on unrelated data, allowing us to distinguish targeted forgetting from broader representational disruption.
(a) ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(b) ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(c) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(e) ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(f) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
Figure 2: Layer-wise PCA Similarity for GA on Yi-6B (simple task). (a–c) vary LR $\{3\times 10^{-6},5\times 10^{-6},3\times 10^{-5}\}$ at $N=100$; (d–f) vary $N\in\{6,50,100\}$ at LR $=3\times 10^{-5}$. Similarity near 1 signals reversible (catastrophic) forgetting; sustained low similarity signals irreversible (catastrophic) forgetting.
(a) ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(b) ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(c) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(e) ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(f) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
Figure 3: PCA Shift for GA on Yi-6B, simple task. (a–c) LR $\{3\times 10^{-6},5\times 10^{-6},3\times 10^{-5}\}$ with $N=100$; (d–f) LR $=3\times 10^{-5}$ with $N\in\{6,50,100\}$. Shift magnitude reflects feature displacement: large, unrecovered shifts indicate severe, irreversible (catastrophic) forgetting, while small or fully recovered shifts indicate mild, reversible (catastrophic) forgetting.
4.2 Representational Results
Principal Component Analysis: Similarity and Shift. Figures 2 and 3 show that, in continual unlearning, larger learning rates or more removal requests drive PCA Similarity sharply downward and leave large, unrecovered PCA Shifts—clear evidence of irreversible catastrophic forgetting. Under gentler hyper-parameters, similarity stays high and shifts remain bounded; relearning then realigns both metrics, signalling reversible catastrophic forgetting. Remarkably, the choice of relearning data and analyzed data—forget, retain, or even unrelated—makes little difference: all three restore the feature geometry, implying the knowledge was suppressed rather than erased (Figure 10 and  14). Due to space limitations, complete figures, single-unlearning results, and additional methods appear in Appendix A.5 and Appendix A.6.
Centered Kernel Alignment Analysis. Figure 4 tracks layer-wise CKA. With gentle unlearning, CKA stays near 1 and relearning restores it completely—typical of reversible catastrophic forgetting. Stronger updates or many requests push deep-layer CKA sharply lower, and alignment cannot be fully recovered, signalling irreversible catastrophic forgetting. The choice of relearning or probe data—forget, retain, or unrelated—barely affects this outcome: once relearning begins, latent structure resurfaces regardless of input (Figure 18). Due to space limitations, complete figures and additional settings appear in Appendix A.5 and Appendix A.6.
(a) ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(b) ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(c) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(e) ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(f) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
Figure 4: CKA for GA on Yi-6B, simple task. (a–c) LR $\{3\times 10^{-6},5\times 10^{-6},3\times 10^{-5}\}$ with $N=100$; (d–f) LR $=3\times 10^{-5}$ with $N\in\{6,50,100\}$. High CKA (near 1) indicates strong subspace alignment and reversible (catastrophic) forgetting, whereas low CKA (near 0) reflects severe representational drift and irreversible (catastrophic) forgetting.
Fisher Information Analysis. Continual unlearning progressively flattens the loss landscape (Figures 5. Higher learning rates or larger request counts push the FIM spectra sharply left, and—at extreme settings—the shift persists after relearning, signalling irreversible catastrophic forgetting. Under gentler hyper-parameters the spectra recentre, indicating reversible catastrophic forgetting. Relearning with forget, retain, or unrelated data realigns the FIM almost equally well, confirming that the lost sensitivity is suppressed rather than erased (Figure 34-38). Due to space limitations, complete figures and additional settings appear in Appendix A.5 and Appendix A.6.
(a) LR=$3\times 10^{-6}$, Layer 31
(b) LR=$5\times 10^{-6}$, Layer 31
(c) LR=$3\times 10^{-5}$, Layer 31
(d) $N=6$, Layer 31
(e) $N=50$, Layer 31
(f) $N=100$, Layer 31
Figure 5: FIM (layer 31) for GA on Yi-6B, simple task. (a–c) LR $\{3\times 10^{-6},5\times 10^{-6},3\times 10^{-5}\}$ with $N=100$; (d–f) LR $=3\times 10^{-5}$ with $N\in\{6,50,100\}$. Larger leftward shifts indicate a greater flattening of the loss landscape and irreversible (catastrophic) forgetting; spectra that recenter on the original peak denote reversible (catastrophic) forgetting.
4.3 Representational Theoretical Analysis
To explain the empirical distinction between reversible and irreversible (catastrophic) forgetting, we present a perturbation model linking unlearning updates to structural collapse across layers. Consider an $L$-layer feedforward neural network $f(x)=\sigma(W_{L}\,\sigma(\cdots\sigma(W_{1}x)\cdots))$, with activations $\sigma$ and weights ${W_{i}}_{i=1}^{L}$. Unlearning is modeled as layer-wise perturbations $\widetilde{W}_{i}=W_{i}+E_{i}$ with $|E_{i}|=O(\mathrm{LR},N)$, where $\mathrm{LR}$ is the learning rate and $N$ the number of unlearning steps. A Neumann-series expansion yields $\widetilde{f}(x)-f(x)=\sum_{\emptyset\neq S\subseteq\{1,\dots,L\}}(W_{L}\circ\cdots\circ E_{i_{k}}\circ\cdots\circ W_{1})(x)$.
When small perturbations are confined to a few layers, first-order terms dominate, and the effect is reversible (catastrophic) forgetting. In contrast, comparable perturbations spread across many layers, higher-order terms accumulate, producing irreversible (catastrophic) forgetting.
PCA Similarity. Let $X_{i}$ and $Y_{i}=X_{i}+E^{\prime}_{i}$ be the centered activations at layer $i$ before and after unlearning. By Davis–Kahan theorem [5], $\cos\angle(\mathbf{c}_{i}^{\text{orig}},\mathbf{c}_{i}^{\text{upd}})\approx 1-O(\|E_{i}^{\prime}\|/(\lambda_{1,i}-\lambda_{2,i})),$ with top two eigenvalues $\lambda_{1,i},\lambda_{2,i}$. The layer-averaged PCA similarity is $\bar{S}_{\mathrm{PCA}}\approx 1-O((1/L)\sum_{i}\|E^{\prime}_{i}\|)$.
PCA Shift. Along the first principal component, the activation-centroid shift is $\Delta p_{i}=\mu_{i,1}^{\text{upd}}-\mu_{i,1}^{\text{orig}}=O(\|E_{i}^{\prime}\|)$. Large perturbations $\|E_{i}^{\prime}\|$ spanning many layers cause irreversible representational drift; otherwise, the shifts remain localized and reversible.
CKA. Let $\widetilde{K}_{Y_{i}}=\widetilde{K}_{X_{i}}+\Delta K_{i}$ be the perturbed Gram matrix. Then, $\mathrm{CKA}_{i}$ is computed as $1-O(\|\Delta K_{i}\|_{*}/\|\widetilde{K}_{X_{i}}\|_{*})$, which implies that $\bar{C}\approx 1-O((1/L)\sum_{i}\|\Delta K_{i}\|_{*})$.
Fisher Information. Given update $\delta w_{i}=O(\|E_{i}\|)$, the Fisher diagonal behaves as $F_{ii}(w+\delta w)=F_{ii}(w)+O(\|\delta w_{i}\|)$, so the average Fisher becomes $\bar{F}=(1/P)\sum_{i}F_{ii}=F_{0}-O((1/P)\sum_{i}\|E_{i}\|)$.
Token-level metrics, such as accuracy, MIA, and AUC, may report total collapse even when the model’s internal geometry is largely preserved. In reversible forgetting, a few parameter changes (e.g., in output heads or layer norms) can drastically perturb token probabilities while leaving deeper representations intact. For the soft-max output $\log p(y|x;\theta+\delta\theta)\approx\log p(y|x;\theta)+\nabla_{\theta}\log p(y|x;\theta)^{\top}\delta\theta+O(\|\delta\theta\|^{2}).$ a small $\delta\theta$ in high-sensitivity regions (near the logits) can dominate the first-order term, producing large drops in accuracy or anomalous AUC scores despite minimal representational drift.
Unlearning on the forget set applies a weight update $E_{i}=LR\times\,\nabla_{W_{i}}L(D_{f}),$ which both removes over-fitting to $\mathcal{D}_{f}$ and accentuates its principal feature subspace. After relearning, parameters may return close to their originals, yet amplified patterns persist, sometimes yielding better performance on an augmented $\mathcal{D}_{f}$ than the baseline. Therefore, unlearning can inadvertently act as a contrastive regularizer, further illustrating the mismatch between surface metrics and feature subspaces.
Summary. Reversible (catastrophic) forgetting occurs when perturbations touch only a few layers; PCA similarity/shift, CKA, and FIM remain near baseline. In contrast, irreversible (catastrophic) forgetting emerges from large, distributed updates that collapse the model’s representational structure.
Unlearning acts as a contrastive perturbation: it removes memorized content yet “reinforces” salient features of the forget set, so subsequent relearning can even outperform the original model on related inputs, revealing its “dual role” in both removal and refinement.
Token-level metrics (e.g., accuracy, MIA) are overly sensitive to small shifts in high-impact parameters and can misclassify the regime. Structural diagnostics, complemented by augmented evaluation, can provide a more reliable assessment of whether forgetting is truly irreversible.
5 Discussion and Takeaways
Beyond theoretical justifications in Section 4, we summarize main empirical and analytical insights.
(1) Single vs. continual unlearning, and the role of GA/RLabel.
Single unlearning rarely produces irreversible collapse: performance is recoverable and representational drift is slight. Continual unlearning, especially with large learning rates, often drives the model into permanent failure: $\sim 100$ sequential requests can push both forget- and retain-set accuracy to near zero. GA and RLabel already over-forget in the single scenario and magnify this damage when applied continually. Adding retain-set terms, as in GA+KL or NPO(+KL) [33; 35; 31], markedly improves stability.
(2) Collapse stems from structural drift, not true erasure.
PCA-Similarity/Shift, CKA, and the FIM consistently expose this breakdown: irreversible collapse coincides with large rotations of principal directions, centroid shifts, and vanishing Fisher mass across many layers. When perturbations remain local (small LR, few unlearning requests), these diagnostics stay near baseline—reversible forgetting. Token-level metrics alone are unreliable: small updates in high-sensitivity parameters (e.g., logits or layer norms) can tank accuracy or inflate MIA AUC while internal geometry is intact.
(3) Unlearning can act as implicit augmentation.
In several continual runs, subsequent relearning on the forget set can often yield higher accuracy than that of the original model. This surprising outcome suggests that unlearning is not merely a memory deletion mechanism but may also serve as a form of implicit contrastive regularization. As detailed in Section 4, unlearning amplifies the feature subspace associated with the forget set, and relearning on augmented inputs can reinforce semantic structure while promoting robustness. This process reorganizes internal representations to better capture generalizable patterns, acting as a form of curriculum learning.
(4) Diagnostics guide irreversible (benign) forgetting.
Tools such as PCA Similarity and Shift, CKA, and Fisher Information reveal not just the presence of structural drift, but also where and how it arises across layers. This enables targeted control: effective unlearning can be guided toward perturbing only those parameters responsible for the forget set, while preserving the representation structure on the retain set and unrelated data. This opens a path to targeted, irreversible forgetting, a permanent and isolated removal of information without collateral collapse, offering actionable insights for building safer unlearning algorithms.
6 Conclusion
We revisit machine unlearning for LLMs through a systematic study of reversibility. Token-level metrics alone can mislead: a model may appear collapsed yet remain fully recoverable.
To diagnose this gap, we introduce a feature-space toolkit—PCA Similarity, PCA Shift, CKA, and FIM—that cleanly separates reversible from irreversible catastrophic forgetting. Our empirical and theoretical results show that true forgetting arises only when many layers undergo coordinated, large-magnitude perturbations; by contrast, minor updates in high-sensitivity regions (e.g., output logits) can slash accuracy or inflate perplexity while leaving internal representations intact.
These findings call for evaluation protocols that go beyond token-level scores and for algorithms that actively control representational drift. We further observe that unlearning, followed by proper relearning, can refine representations and even boost downstream performance. Together, these insights chart a path toward safer, more interpretable unlearning in LLMs.
Appendix A Appendix
A.1 Limitations
Our experiments target two LLMs and a handful of tasks and unlearning methods; although our diagnostic framework is model-agnostic and designed to scale, empirical validation on much larger models and production-scale pipelines remains to be done. The constrained relearning protocol and selected metrics provide clear insights into representational drift but are not exhaustive and do not offer formal privacy guarantees. Extending these analyses to diverse architectures, richer unlearning strategies, and rigorous theoretical bounds is an important direction for future work.
A.2 Broader impacts
This work provides new tools and theoretical insights for evaluating and guiding machine unlearning in LLMs, with the potential to improve privacy guarantees and regulatory compliance (e.g., GDPR “right to be forgotten”). By distinguishing reversible from irreversible forgetting, our diagnostics can help developers ensure that sensitive or copyrighted material is truly expunged rather than merely hidden. At the same time, adversaries could exploit reversible forgetting pathways to mask malicious or biased content and then restore it later, highlighting the need for robust defenses. The computational overhead of continual unlearning and repeated diagnostics may increase energy consumption, underscoring the importance of efficient implementations. Finally, by exposing the fragility of token-level metrics, our work advocates for more trustworthy evaluation standards that balance privacy, utility, and environmental considerations in real-world deployments.
A.3 Detailed Analysis Tools
PCA Similarity and PCA Shift.
For each Transformer layer, we perform PCA on the hidden activations of the original and updated models. Let $\mathbf{c}_{i,1}^{\text{orig}}$ and $\mathbf{c}_{i,1}^{\text{upd}}$ denote the first principal component (PC1) directions of layer $i$. The PCA Similarity is defined as
$\operatorname{PCA\text{-}Sim}(i)=\cos\bigl(\mathbf{c}_{i,1}^{\text{orig}},\mathbf{c}_{i,1}^{\text{upd}}\bigr{missing})=\frac{(\mathbf{c}_{i,1}^{\text{orig}})^{\top}\mathbf{c}_{i,1}^{\text{upd}}}{\lVert\mathbf{c}_{i,1}^{\text{orig}}\rVert\,\lVert\mathbf{c}_{i,1}^{\text{upd}}\rVert}\in[-1,1],$
where values near $1$ indicate stable directional alignment, and values near $-1$ suggest a near-orthogonal shift in dominant directions.
To capture translational drift, we also compute the mean projection of activations along PC1 and PC2:
$\operatorname{PCA\text{-}Shift}(i)=p_{1,\text{upd}}-p_{1,\text{orig}},\qquad\operatorname{Principle}(i)=p_{2,\text{upd}},$
where $\operatorname{PCA\text{-}Shift}$ quantifies displacement along PC1 and $\operatorname{Principle}$ captures orthogonal deviation along PC2. These metrics reflect how the representation center drifts within the top subspace.
Centered Kernel Alignment (CKA).
To assess subspace alignment, we use linear Centered Kernel Alignment (CKA) [15], which compares activation matrices $X,Y\in\mathbb{R}^{N\times D}$ from before and after unlearning. First, we compute the centered Gram matrices:
$\widetilde{K}_{X}=HXX^{\top}H,\qquad\widetilde{K}_{Y}=HYY^{\top}H,\qquad H=I_{N}-\tfrac{1}{N}\mathbf{1}\mathbf{1}^{\top}.$
The CKA score is then given by:
$\operatorname{CKA}(X,Y)=\frac{\operatorname{Tr}(\widetilde{K}_{X}\widetilde{K}_{Y})}{\sqrt{\operatorname{Tr}(\widetilde{K}_{X}^{2})}\sqrt{\operatorname{Tr}(\widetilde{K}_{Y}^{2})}}\in[0,1],$
where values near $1$ indicate highly overlapping subspaces, and values near 00 signal near-orthogonality.
Fisher Information.
To measure parameter-level importance, we compute the diagonal of the empirical Fisher Information Matrix (FIM). For each parameter $w_{i}$ and input distribution $\mathcal{D}_{\text{dis}}$, the diagonal entry is approximated as:
$\operatorname{FIM}_{ii}\approx\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}_{\text{dis}}}\left[\left(\partial_{w_{i}}\log p(y\mid\mathbf{x};\mathbf{w})\right)^{2}\right].$
Larger values indicate that $w_{i}$ has a stronger influence on the model’s predictions. A substantial leftward shift in the Fisher spectrum after unlearning implies a flattened loss landscape and diminished parameter sensitivity.
Together, these tools form a feature-space diagnostic suite: FIM captures global sensitivity, CKA measures subspace preservation, and PCA-based metrics expose fine-grained geometric drift across layers—enabling a robust assessment of representational degradation during unlearning.
A.4 Complete Experiment
Table 4: Yi-6B simple-task metrics under four $(\mathrm{LR},N)$ settings. For each block: forget/retain perplexity (F.Ppl / R.Ppl), forget/retain accuracy (F.Acc / R.Acc), CommonsenseQA (CSQA), GSM8K, and membership-inference AUC (MIA).
Phase Method F.Ppl R.Ppl F.Acc R.Acc CSQA GSM8K MIA
LR=3×$1003\times 10^{-5},\,N=$ 100
Original 3.8 7.8 78.9 65.5 73.1 39.6 70.9
Unlearn GA $\infty$ $\infty$ 0.0 0.0 19.3 0.0 26.1
GA+GD $\infty$ $\infty$ 9.7 2.3 19.7 0.0 16.8
GA+KL $\infty$ $\infty$ 9.0 6.2 19.6 0.0 17.8
NPO 31296.5 597.9 37.8 37.9 62.2 1.0 60.1
NPO+KL 348080.2 4482.0 64.3 55.9 64.9 1.4 59.0
Rlable 63791.7 65903.4 0.0 0.0 20.9 0.0 65.1
Relearn GA 137094.5 758443.5 2.1 1.8 19.7 0.0 74.5
GA+GD 5274.5 9568.6 2.2 2.6 19.6 0.0 68.1
GA+KL 5037.1 15019.9 1.7 1.6 20.6 0.0 70.7
NPO 16.6 41.7 57.0 45.6 51.8 0.6 70.0
NPO+KL 21.8 16.2 60.7 54.3 48.0 0.9 67.7
Rlable 4056.1 15048.6 4.3 2.8 19.7 0.0 69.5
LR=5×$1005\times 10^{-6},\,N=$ 100
Unlearn GA $\infty$ $\infty$ 9.1 6.2 19.6 0.0 23.2
GA+GD $\infty$ $\infty$ 3.6 3.1 24.5 0.0 28.7
GA+KL $\infty$ $\infty$ 9.1 6.2 19.6 0.0 27.3
NPO 3017.7 1110.6 50.1 52.3 72.9 37.5 50.6
NPO+KL 38.5 232.4 77.6 64.3 73.1 37.6 65.4
Rlable 57035.4 53377.1 0.1 0.4 19.1 0.0 63.6
Relearn GA 3.7 7.8 80.0 64.9 70.2 39.9 68.0
GA+GD 3.6 7.6 81.2 65.1 72.1 39.0 69.8
GA+KL 3.6 8.4 81.1 64.8 71.6 40.7 68.3
NPO 3.5 7.6 82.7 65.5 74.0 39.7 68.0
NPO+KL 3.5 7.8 83.8 65.6 74.1 39.7 69.5
Rlable 3.6 7.7 80.8 65.3 71.8 39.2 70.3
LR=3×$1003\times 10^{-6},\,N=$ 100
Unlearn GA $\infty$ $\infty$ 16.8 14.4 69.5 12.3 25.2
GA+GD 3.3 7.6 78.8 65.5 77.0 37.5 69.4
GA+KL $\infty$ $\infty$ 35.4 40.6 63.2 18.3 18.9
NPO 3.7 7.9 78.3 65.0 73.3 38.7 68.4
NPO+KL 3.8 8.1 78.4 65.1 73.6 38.6 66.7
Rlable 36794.7 32562.0 3.8 3.2 19.3 2.2 61.4
Relearn GA 3.7 7.6 80.8 65.2 73.4 39.9 68.6
GA+GD 3.6 7.4 81.8 65.5 72.1 39.0 70.0
GA+KL 3.6 10.3 81.0 63.3 67.2 40.7 70.7
NPO 3.5 7.5 81.2 65.4 72.9 39.7 69.9
NPO+KL 3.5 7.5 83.8 65.5 73.0 39.7 69.9
Rlable 3.6 7.6 80.5 65.3 72.2 39.2 70.0
LR=$3\times 10^{-5},\,N=6$
Unlearn GA inf inf 36.3 36.1 69.1 5.8 29.6
GA+GD 209.3 20.6 77.0 64.0 70.0 37.8 66.9
GA+KL inf inf 53.0 41.5 68.3 2.0 29.5
NPO 12.3 10.7 71.6 59.4 71.7 24.7 68.7
NPO+KL 8.9 10.7 74.7 62.1 72.8 32.2 67.9
Rlable 51589.2 40622.9 0.4 0.7 19.8 0.0 62.6
Relearn GA 6.8 11.4 70.5 58.7 64.5 18.4 68.2
GA+GD 12.3 11.5 61.6 54.4 61.3 7.3 67.1
GA+KL 17.1 11.6 66.6 56.2 60.6 3.0 65.0
NPO 6.0 11.6 71.2 59.4 59.4 2.0 68.4
NPO+KL 7.3 11.6 67.6 56.1 42.9 1.6 69.0
Rlable 6.4 11.4 72.7 61.1 67.5 28.9 65.2
Table 4 summarizes token-level performance under four unlearning configurations on Yi-6B, varying both learning rates and request counts. Most metrics—including forget/retain accuracy, perplexity, CSQA, and GSM8K—follow consistent trends: milder methods (e.g., NPO, NPO+KL) yield moderate degradation and support effective recovery via relearning, while aggressive methods (e.g., GA, RL) under high learning rates lead to severe, often irrecoverable performance drops. These results align closely with our theoretical distinction: reversible (catastrophic) forgetting corresponds to localized perturbations and structurally recoverable models, while irreversible (catastrophic) forgetting reflects widespread, irreversible damage.
MIA, however, behaves differently. Even when downstream performance collapses, MIA scores often remain high and recover rapidly. This suggests MIA is especially sensitive to shallow changes (e.g., output logits or normalization), rather than deeper representational shifts. As a result, MIA often correlates with reversible forgetting—capturing surface-level instability without revealing structural collapse. These observations reinforce that while token-level metrics offer partial signals, only internal diagnostics can reliably differentiate reversible from irreversible forgetting.
A.5 Single Unlearning
Figure 6 illustrates feature-level changes under single unlearning using PCA Similarity, PCA Shift, CKA, and Fisher Information. In (a), PCA Similarity remains consistently high across layers, with cosine scores near 1, indicating that dominant activation directions are well preserved. Slight dips in shallow and final layers are quickly restored after relearning, suggesting minimal and reversible geometric drift. Subfigure (b) confirms that PC1 shifts and orthogonal deviations are small, with relearned centers closely matching the original. In (c), CKA shows near-perfect alignment between the original, unlearned, and relearned representations, reinforcing the conclusion that subspace structure remains intact. Fisher Information spectra in (d–f) reveal only mild leftward shifts, indicating slight loss flattening and reduced parameter sensitivity, which are fully recovered after relearning. Overall, these results confirm that single unlearning causes only minor, reversible structural perturbations—highlighting the fragility of token-level evaluations in capturing irreversible forgetting.
(a) Simple (Single Unlearning)
(b) Simple (Single Unlearning)
(c) Simple (Single Unlearning)
(d) Simple (Single Unlearning)
(e) Simple (Single Unlearning)
(f) Simple (Single Unlearning)
Figure 6: Single unlearning analysis on Yi-6B for GA under a simple task. PCA Similarity, PCA Shift, CKA, and Fisher information (FIM) distributions are reported across key layers to assess representation and parameter changes before and after unlearning and relearning.
A.6 Detailed Analysis Results
(a) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(b) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(c) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(e) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(f) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(g) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(h) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(i) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(j) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(k) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(l) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(m) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(n) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(o) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
[…]
(a) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(b) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(c) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(e) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(f) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(g) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(h) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(i) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(j) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(k) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(l) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(m) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(n) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(o) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
Figure 8: PCA Similarity Across Layers. Each row shows results under different unlearning methods: GA+GD (a–c), GA+KL (d–f), NPO (g–i), NPO+KL (j–l), and Rlable (m–o). Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
(a) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
(b) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
(c) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(d) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
(e) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
(f) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(g) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
(h) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
(i) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
[…]
(a) Simple (Relearned by forget set)
(b) Simple (Relearned by Retain set)
(c) Simple (Relearned by unrelated data)
(d) Simple (input data = forget set)
(e) Simple (input data = retain set)
(f) Simple (input data = unrelated data)
Figure 10: PCA Similarity Analysis for GA under Varied Relearning and Evaluation Inputs on Yi-6B (Simple Task). (a–c): Relearning is performed using the forget set, retain set, or unrelated data respectively. (d–f): PCA similarity is measured using the forget set, retain set, or unrelated data as evaluation input.
(a) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(b) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(c) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(e) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(f) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(g) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(h) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(i) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(j) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(k) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(l) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(m) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(n) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(o) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
[…]
(a) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(b) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(c) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(e) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(f) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(g) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(h) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(i) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(j) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(k) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(l) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(m) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(n) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(o) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
Figure 12: PCA Shift Across Layers. Each row shows results under different unlearning methods: GA+GD (a–c), GA+KL (d–f), NPO (g–i), NPO+KL (j–l), and Rlable (m–o). Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
(a) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(b) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(c) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(e) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(f) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(g) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(h) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(i) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
[…]
(a) Simple (Relearned by forget set)
(b) Simple (Relearned by retain set)
(c) Simple (Relearned by unrelated data)
(d) Simple (input data = forget set)
(e) Simple (input data = retain set)
(f) Simple (input data = unrelated data)
Figure 14: PCA Shift Analysis under Varied Relearning and Evaluation Inputs on Yi-6B (Simple Task). (a–c): Relearning is performed using the forget set, retain set, or unrelated data respectively. (d–f): PCA shift is measured using the forget set, retain set, or unrelated data as evaluation input.
(a) Simple Task ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(b) Simple Task ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(c) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) Simple Task ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(e) Simple Task ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(f) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(g) Simple Task ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(h) Simple Task ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(i) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(j) Simple Task ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(k) Simple Task ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(l) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(m) Simple Task ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(n) Simple Task ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(o) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
[…]
(a) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(b) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(c) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(e) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(f) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(g) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(h) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(i) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(j) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(k) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(l) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(m) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(n) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(o) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
Figure 16: CKA Across Layers. Each row shows results under different unlearning methods: GA+GD (a–c), GA+KL (d–f), NPO (g–i), NPO+KL (j–l), and Rlable (m–o). Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
(a) Complex Task ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
(b) Complex Task ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
(c) Complex Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(d) Complex Task ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
(e) Complex Task ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
(f) Complex Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(g) Complex Task ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
(h) Complex Task ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
(i) Complex Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
[…]
(a) Simple (Relearned by forget set)
(b) Simple (Relearned by retain set)
(c) Simple (Relearned by unrelated data)
(d) Simple (input data = forget set)
(e) Simple (input data = retain set)
(f) Simple (input data = unrelated data)
Figure 18: CKA Analysis under Varied Relearning and Evaluation Inputs on Yi-6B (Simple Task). (a–c): Relearning is performed using the forget set, retain set, or unrelated data respectively. (d–f): CKA is measured using the forget set, retain set, or unrelated data as evaluation input.
(a) LR=$3\times 10^{-6}$, Layer 28
(b) LR=$5\times 10^{-6}$, Layer 28
(c) LR=$3\times 10^{-5}$, Layer 28
(d) LR=$3\times 10^{-6}$, Layer 22
(e) LR=$5\times 10^{-6}$, Layer 22
(f) LR=$3\times 10^{-5}$, Layer 22
(g) LR=$3\times 10^{-6}$, Layer 13
(h) LR=$5\times 10^{-6}$, Layer 13
(i) LR=$3\times 10^{-5}$, Layer 13
(j) LR=$3\times 10^{-6}$, Layer 4
(k) LR=$5\times 10^{-6}$, Layer 4
(l) LR=$3\times 10^{-5}$, Layer 4
(m) LR=$3\times 10^{-6}$, Layer 1
(n) LR=$5\times 10^{-6}$, Layer 1
(o) LR=$3\times 10^{-5}$, Layer 1
[…]
(a) LR=$3\times 10^{-6}$, Layer 31
(b) LR=$5\times 10^{-6}$, Layer 31
(c) LR=$3\times 10^{-5}$, Layer 31
(d) LR=$3\times 10^{-6}$, Layer 28
(e) LR=$5\times 10^{-6}$, Layer 28
(f) LR=$3\times 10^{-5}$, Layer 28
(g) LR=$3\times 10^{-6}$, Layer 22
(h) LR=$5\times 10^{-6}$, Layer 22
(i) LR=$3\times 10^{-5}$, Layer 22
(j) LR=$3\times 10^{-6}$, Layer 13
(k) LR=$5\times 10^{-6}$, Layer 13
(l) LR=$3\times 10^{-5}$, Layer 13
(m) LR=$3\times 10^{-6}$, Layer 4
(n) LR=$5\times 10^{-6}$, Layer 4
(o) LR=$3\times 10^{-5}$, Layer 4
(p) LR=$3\times 10^{-6}$, Layer 1
(q) LR=$5\times 10^{-6}$, Layer 1
(r) LR=$3\times 10^{-5}$, Layer 1
[…]
(a) LR=$3\times 10^{-6}$, Layer 31
(b) LR=$5\times 10^{-6}$, Layer 31
(c) LR=$3\times 10^{-5}$, Layer 31
(d) LR=$3\times 10^{-6}$, Layer 28
(e) LR=$5\times 10^{-6}$, Layer 28
(f) LR=$3\times 10^{-5}$, Layer 28
(g) LR=$3\times 10^{-6}$, Layer 22
(h) LR=$5\times 10^{-6}$, Layer 22
(i) LR=$3\times 10^{-5}$, Layer 22
(j) LR=$3\times 10^{-6}$, Layer 13
(k) LR=$5\times 10^{-6}$, Layer 13
(l) LR=$3\times 10^{-5}$, Layer 13
(m) LR=$3\times 10^{-6}$, Layer 4
(n) LR=$5\times 10^{-6}$, Layer 4
(o) LR=$3\times 10^{-5}$, Layer 4
(p) LR=$3\times 10^{-6}$, Layer 1
(q) LR=$5\times 10^{-6}$, Layer 1
(r) LR=$3\times 10^{-5}$, Layer 1
[…]
(a) LR=$3\times 10^{-6}$, Layer 31
(b) LR=$5\times 10^{-6}$, Layer 31
(c) LR=$3\times 10^{-5}$, Layer 31
(d) LR=$3\times 10^{-6}$, Layer 28
(e) LR=$5\times 10^{-6}$, Layer 28
(f) LR=$3\times 10^{-5}$, Layer 28
(g) LR=$3\times 10^{-6}$, Layer 22
(h) LR=$5\times 10^{-6}$, Layer 22
(i) LR=$3\times 10^{-5}$, Layer 22
(j) LR=$3\times 10^{-6}$, Layer 13
(k) LR=$5\times 10^{-6}$, Layer 13
(l) LR=$3\times 10^{-5}$, Layer 13
(m) LR=$3\times 10^{-6}$, Layer 4
(n) LR=$5\times 10^{-6}$, Layer 4
(o) LR=$3\times 10^{-5}$, Layer 4
(p) LR=$3\times 10^{-6}$, Layer 1
(q) LR=$5\times 10^{-6}$, Layer 1
(r) LR=$3\times 10^{-5}$, Layer 1
[…]
(a) LR=$3\times 10^{-6}$, Layer 31
(b) LR=$5\times 10^{-6}$, Layer 31
(c) LR=$3\times 10^{-5}$, Layer 31
(d) LR=$3\times 10^{-6}$, Layer 28
(e) LR=$5\times 10^{-6}$, Layer 28
(f) LR=$3\times 10^{-5}$, Layer 28
(g) LR=$3\times 10^{-6}$, Layer 22
(h) LR=$5\times 10^{-6}$, Layer 22
(i) LR=$3\times 10^{-5}$, Layer 22
(j) LR=$3\times 10^{-6}$, Layer 13
(k) LR=$5\times 10^{-6}$, Layer 13
(l) LR=$3\times 10^{-5}$, Layer 13
(m) LR=$3\times 10^{-6}$, Layer 4
(n) LR=$5\times 10^{-6}$, Layer 4
(o) LR=$3\times 10^{-5}$, Layer 4
(p) LR=$3\times 10^{-6}$, Layer 1
(q) LR=$5\times 10^{-6}$, Layer 1
(r) LR=$3\times 10^{-5}$, Layer 1
[…]
(a) LR=$3\times 10^{-6}$, Layer 31
(b) LR=$5\times 10^{-6}$, Layer 31
(c) LR=$3\times 10^{-5}$, Layer 31
(d) LR=$3\times 10^{-6}$, Layer 28
(e) LR=$5\times 10^{-6}$, Layer 28
(f) LR=$3\times 10^{-5}$, Layer 28
(g) LR=$3\times 10^{-6}$, Layer 22
(h) LR=$5\times 10^{-6}$, Layer 22
(i) LR=$3\times 10^{-5}$, Layer 22
(j) LR=$3\times 10^{-6}$, Layer 13
(k) LR=$5\times 10^{-6}$, Layer 13
(l) LR=$3\times 10^{-5}$, Layer 13
(m) LR=$3\times 10^{-6}$, Layer 4
(n) LR=$5\times 10^{-6}$, Layer 4
(o) LR=$3\times 10^{-5}$, Layer 4
(p) LR=$3\times 10^{-6}$, Layer 1
(q) LR=$5\times 10^{-6}$, Layer 1
(r) LR=$3\times 10^{-5}$, Layer 1
[…]
(a) $N=6$, Layer 28
(b) $N=50$, Layer 28
(c) $N=100$, Layer 28
(d) $N=6$, Layer 22
(e) $N=50$, Layer 22
(f) $N=100$, Layer 22
(g) $N=6$, Layer 13
(h) $N=50$, Layer 13
(i) $N=100$, Layer 13
(j) $N=6$, Layer 4
(k) $N=50$, Layer 4
(l) $N=100$, Layer 4
(m) $N=6$, Layer 1
(n) $N=50$, Layer 1
(o) $N=100$, Layer 1
Figure 25: FIM for GA Across Layers. Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
(a) $N=6$, Layer 31
(b) $N=50$, Layer 31
(c) $N=100$, Layer 31
(d) $N=6$, Layer 28
(e) $N=50$, Layer 28
(f) $N=100$, Layer 28
(g) $N=6$, Layer 22
(h) $N=50$, Layer 22
(i) $N=100$, Layer 22
(j) $N=6$, Layer 13
(k) $N=50$, Layer 13
(l) $N=100$, Layer 13
(m) $N=6$, Layer 4
(n) $N=50$, Layer 4
(o) $N=100$, Layer 4
(p) $N=6$, Layer 1
(q) $N=50$, Layer 1
(r) $N=100$, Layer 1
Figure 26: FIM for GA+GD Across Layers. Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
(a) $N=6$, Layer 31
(b) $N=50$, Layer 31
(c) $N=100$, Layer 31
(d) $N=6$, Layer 28
(e) $N=50$, Layer 28
(f) $N=100$, Layer 28
(g) $N=6$, Layer 22
(h) $N=50$, Layer 22
(i) $N=100$, Layer 22
(j) $N=6$, Layer 13
(k) $N=50$, Layer 13
(l) $N=100$, Layer 13
(m) $N=6$, Layer 4
(n) $N=50$, Layer 4
(o) $N=100$, Layer 4
(p) $N=6$, Layer 1
(q) $N=50$, Layer 1
(r) $N=100$, Layer 1
Figure 27: FIM for GA+KL Across Layers. Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
(a) $N=6$, Layer 31
(b) $N=50$, Layer 31
(c) $N=100$, Layer 31
(d) $N=6$, Layer 28
(e) $N=50$, Layer 28
(f) $N=100$, Layer 28
(g) $N=6$, Layer 22
(h) $N=50$, Layer 22
(i) $N=100$, Layer 22
(j) $N=6$, Layer 13
(k) $N=50$, Layer 13
(l) $N=100$, Layer 13
(m) $N=6$, Layer 4
(n) $N=50$, Layer 4
(o) $N=100$, Layer 4
(p) $N=6$, Layer 1
(q) $N=50$, Layer 1
(r) $N=100$, Layer 1
Figure 28: FIM for NPO Across Layers. Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
(a) $N=6$, Layer 31
(b) $N=50$, Layer 31
(c) $N=100$, Layer 31
(d) $N=6$, Layer 28
(e) $N=50$, Layer 28
(f) $N=100$, Layer 28
(g) $N=6$, Layer 22
(h) $N=50$, Layer 22
(i) $N=100$, Layer 22
(j) $N=6$, Layer 13
(k) $N=50$, Layer 13
(l) $N=100$, Layer 13
(m) $N=6$, Layer 4
(n) $N=50$, Layer 4
(o) $N=100$, Layer 4
(p) $N=6$, Layer 1
(q) $N=50$, Layer 1
(r) $N=100$, Layer 1
Figure 29: FIM for NPO+KL Across Layers. Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
(a) $N=6$, Layer 31
(b) $N=50$, Layer 31
(c) $N=100$, Layer 31
(d) $N=6$, Layer 28
(e) $N=50$, Layer 28
(f) $N=100$, Layer 28
(g) $N=6$, Layer 22
(h) $N=50$, Layer 22
(i) $N=100$, Layer 22
(j) $N=6$, Layer 13
(k) $N=50$, Layer 13
(l) $N=100$, Layer 13
(m) $N=6$, Layer 4
(n) $N=50$, Layer 4
(o) $N=100$, Layer 4
(p) $N=6$, Layer 1
(q) $N=50$, Layer 1
(r) $N=100$, Layer 1
Figure 30: FIM for Rlable Across Layers. Simple task on Yi-6B with fixed learning rate $\mathrm{LR}=3\times 10^{-5}$ and varying unlearning requests $N\in\{6,50,100\}$.
(a) LR=$3\times 10^{-6}$, Layer 28
(b) LR=$5\times 10^{-6}$, Layer 28
(c) LR=$3\times 10^{-5}$, Layer 28
(d) LR=$3\times 10^{-6}$, Layer 24
(e) LR=$5\times 10^{-6}$, Layer 24
(f) LR=$3\times 10^{-5}$, Layer 24
(g) LR=$3\times 10^{-6}$, Layer 12
(h) LR=$5\times 10^{-6}$, Layer 12
(i) LR=$3\times 10^{-5}$, Layer 12
(j) LR=$3\times 10^{-6}$, Layer 4
(k) LR=$5\times 10^{-6}$, Layer 4
(l) LR=$3\times 10^{-5}$, Layer 4
(m) LR=$3\times 10^{-6}$, Layer 1
(n) LR=$5\times 10^{-6}$, Layer 1
(o) LR=$3\times 10^{-5}$, Layer 1
[…]
(a) LR=$3\times 10^{-6}$, Layer 28
(b) LR=$5\times 10^{-6}$, Layer 28
(c) LR=$3\times 10^{-5}$, Layer 28
(d) LR=$3\times 10^{-6}$, Layer 24
(e) LR=$5\times 10^{-6}$, Layer 24
(f) LR=$3\times 10^{-5}$, Layer 24
(g) LR=$3\times 10^{-6}$, Layer 12
(h) LR=$5\times 10^{-6}$, Layer 12
(i) LR=$3\times 10^{-5}$, Layer 12
(j) LR=$3\times 10^{-6}$, Layer 4
(k) LR=$5\times 10^{-6}$, Layer 4
(l) LR=$3\times 10^{-5}$, Layer 4
(m) LR=$3\times 10^{-6}$, Layer 1
(n) LR=$5\times 10^{-6}$, Layer 1
(o) LR=$3\times 10^{-5}$, Layer 1
[…]
(a) LR=$3\times 10^{-6}$, Layer 28
(b) LR=$5\times 10^{-6}$, Layer 28
(c) LR=$3\times 10^{-5}$, Layer 28
(d) LR=$3\times 10^{-6}$, Layer 24
(e) LR=$5\times 10^{-6}$, Layer 24
(f) LR=$3\times 10^{-5}$, Layer 24
(g) LR=$3\times 10^{-6}$, Layer 12
(h) LR=$5\times 10^{-6}$, Layer 12
(i) LR=$3\times 10^{-5}$, Layer 12
(j) LR=$3\times 10^{-6}$, Layer 4
(k) LR=$5\times 10^{-6}$, Layer 4
(l) LR=$3\times 10^{-5}$, Layer 4
(m) LR=$3\times 10^{-6}$, Layer 1
(n) LR=$5\times 10^{-6}$, Layer 1
(o) LR=$3\times 10^{-5}$, Layer 1
[…]
(a) Simple (Relearned by forget set), Layer 31
(b) Simple (Relearned by retain set), Layer 31
(c) Simple (Relearned by unrelated data), Layer 31
(d) Simple (input data = forget set), Layer 31
(e) Simple (input data = retain set), Layer 31
(f) Simple (input data = unrelated data), Layer 31
Figure 34: FIM in layer 31 under Varied Relearning and Evaluation Inputs on Yi-6B (Simple Task). (a–c): Relearning is performed using the forget set, retain set, or unrelated data respectively. (d–f): FIM is measured using the forget set, retain set, or unrelated data as evaluation input.
(a) Simple (Relearned by forget set), Layer 25
(b) Simple (Relearned by retain set), Layer 25
(c) Simple (Relearned by unrelated data), Layer 25
(d) Simple (input data = forget set), Layer 25
(e) Simple (input data = retain set), Layer 25
(f) Simple (input data = unrelated data), Layer 25
Figure 35: FIM in layer 25 under Varied Relearning and Evaluation Inputs on Yi-6B (Simple Task). (a–c): Relearning is performed using the forget set, retain set, or unrelated data respectively. (d–f): FIM is measured using the forget set, retain set, or unrelated data as evaluation input.
(a) Simple (Relearned by forget set), Layer 16
(b) Simple (Relearned by retain set), Layer 16
(c) Simple (Relearned by unrelated data), Layer 16
(d) Simple (input data = forget set), Layer 16
(e) Simple (input data = retain set), Layer 16
(f) Simple (input data = unrelated data), Layer 16
Figure 36: FIM in layer 16 under Varied Relearning and Evaluation Inputs on Yi-6B (Simple Task). (a–c): Relearning is performed using the forget set, retain set, or unrelated data respectively. (d–f): FIM is measured using the forget set, retain set, or unrelated data as evaluation input.
(a) Simple (Relearned by forget set), Layer 4
(b) Simple (Relearned by retain set), Layer 4
(c) Simple (Relearned by unrelated data), Layer 4
(d) Simple (input data = forget set Layer), 4
(e) Simple (input data = retain set Layer), 4
(f) Simple (input data = unrelated data), Layer 4
Figure 37: FIM in layer 4 under Varied Relearning and Evaluation Inputs on Yi-6B (Simple Task). (a–c): Relearning is performed using the forget set, retain set, or unrelated data respectively. (d–f): FIM is measured using the forget set, retain set, or unrelated data as evaluation input.
(a) Simple (Relearned by forget set), Layer 1
(b) Simple (Relearned by retain set), Layer 1
(c) Simple (Relearned by unrelated data), Layer 1
(d) Simple (input data = forget set), Layer 1
(e) Simple (input data = retain set), Layer 1
(f) Simple (input data = unrelated data), Layer 1
Figure 38: FIM in layer 1 under Varied Relearning and Evaluation Inputs on Yi-6B (Simple Task). (a–c): Relearning is performed using the forget set, retain set, or unrelated data respectively. (d–f): FIM is measured using the forget set, retain set, or unrelated data as evaluation input.
A.6.1 Principal Component Analysis: Similarity and Shift
Across the same hyper-parameter grid, Figures 7– 13 plot the PCA–Shift trajectories—layer-wise displacements of activation centroids along the first two principal directions. For GA-based objectives the pattern mirrors their Similarity curves: as LR rises the orange triangles (unlearned) shoot far from the blue circles (original), especially in deeper layers, and the green squares (relearned) fail to return, producing long grey rays that diagnose irreversible drift. With GA+GD the spread is smaller but still widens sharply at $3{\times}10^{-5}$, confirming that doubling the loss term does not prevent global collapse.
NPO and NPO+KL behave differently. Even at aggressive LR the shifts remain tightly clustered—most layers move $<\!100$ units on pc1 and almost none on pc2—and green squares consistently fall back a little onto the original line segment. RLabel shows an intermediate picture: early layers barely move, while late layers fan out as LR or $N$ grow; nonetheless the rays shorten markedly after relearning, indicating that most distortion is still recoverable.
Task complexity amplifies divergence. On Qwen-2.5-7B the GA rays explode more quickly (Figure 13c,f,i), spanning thousands of pc1 units and driving pc2 to extreme negative values—the hallmark of a high-order, multi-layer perturbation predicted by our theory in Section 4. Conversely, NPO’s clusters expand only modestly, even under the same LR, and contract once relearning is applied.
Taken together, PCA-Shift complements PCA-Similarity: Similarity captures angular mis-alignment, while Shift quantifies translational drift. Their joint reading confirms that GA (with or without GD/KL) is prone to large, irreversible representational displacements, whereas NPO variants and, to a lesser extent, RLabel confine shifts to a regime that remains correctable—consistent with the reversible versus irreversible forgetting boundary observed in our utility experiments.
A.6.2 Centered Kernel Alignment Analysis
Figures 15– 17 report layer-wise linear CKA between the original model and its unlearned / relearned counterparts. Across both Yi-6B and Qwen-2.5-7B, GA again stands out: as LR or $N$ grows its CKA curve close to zero in the final third of the network and never returns, revealing a deep sub-space fracture that matches the irreversible PCA trends. GA+GD and GA+KL modestly attenuate this dip, but still fail to restore full alignment after relearning.
By contrast, NPO and NPO+KL keep CKA higher GA through almost all layers—even under $3{\times}10^{-5}$ or $N{=}100$—and relearning lifts the few layers back to baseline, confirming their perturbations are lower than GA series. RLabel occupies an intermediate position: as LR or N increases, its CKA curve drops rapidly—mirroring GA’s behavior—and ultimately exhibits irreversible forgetting.
Task complexity does not change the ordering but amplifies the gaps: on the math-heavy Qwen benchmark GA’s tail layers fall to almost zero at high LR, whereas NPO keeps higher than GA. Taken together with the PCA-Shift results, CKA shows that only GA-style objectives consistently destroy the encoder–decoder sub-space, while NPO families maintain higher stability than GA series and RLabel induces a moderate, recoverable tilt.
A.6.3 Fisher Information Analysis
[…]